\begin{thebibliography}{77}

% \bibitem{Bab} 
% K. I. Babenko, 
% {\it About the approximation of periodic functions of many variable trigonometric polynomials},
% Doklady Akademii Nauk SSSR, 32, 247--250, 1960.

% \bibitem{BSS} J. Batson, D.A. Spielman, and N. Srivastava, {\it Twice-Ramanujan Sparsifiers}, SIAM Review, 56, 315-334, 2014.

% \bibitem{BSS} J.D. Batson, D.A. Spielman, N. Srivastava, 
% {\it Twice-Ramanujan sparsifiers}, in: STOC’09—Proceedings of the 2009 ACM International Symposium on Theory of Computing, pp. 255–-262, ACM, New York, 2009.

% \bibitem{BKPSU2024}
% F. Bartel, L. K\"ammerer, K. Pozharska, M. Sch\"afer, T. Ullrich,
% {\it
% Exact discretization, tight frames and recovery via D-optimal designs},
% arXiv:2412.02489, 2024.



% \bibitem{BSU} F. Bartel, M. Sch\"afer, and T. Ullrich, {\it Constructive subsampling of finite frames with applications in optimal function recovery}, Applied. Comput. Harmon. Anal. 65, 209--248, 2023.

% \bibitem{Bernardi_Maday1992}
% C. Bernardi and Y. Maday, 
% {\it  Polynomial interpolation results in {S}obolev spaces},
% {J.~Comput. Appl. Math.},
%  {43(1)}, 53--80,
%  {1992},
%     {https://doi.org/10.1016/0377-0427(92)90259-Z}.

% \bibitem{BhJa09}
% R.~Bhatia, T.~Jain, 
% {\it Higher order derivatives and perturbation bounds for determinants},
% Linear Algebra and its Applications 431 (2009), Issue 11, pp. 2102--2108.


% \bibitem{Binev_Bonito_2023}
% P. Binev, A. Bonito, A. Cohen, W. Dahmen, R. DeVore, and G. Petrova,
% {\it Solving PDEs with Incomplete Information}, arXiv:2301.05540. 

\bibitem{Birman_Solomyak}
M. Sh. Birman, M. Z. Solomyak, 
{\it Piecewise-polynomial approximations of
functions of the classes $W^\alpha_p$},
Mathematics of the USSR-Sbornik, Volume 2, Issue 3, 295--317, 1967.
https://doi.org/10.1070/SM1967v002n03ABEH002343

% \bibitem{Bos18}
% L.~Bos, 
% \textit{Fekete points as norming sets}, 
% Dolomites Research Notes on Approximation 11(4), 26--34, 2018. 
% %doi: 10.14658/pupj-drna-2018-4-3

% \bibitem{Bos22}
% %\kp{
% L.~Bos, 
% \textit{On optimal designs for a $d$-cube},
% Dolomites Research Notes on Approximation 15(4), 20--34, 2022.
% %}
% %DOI: 10.14658/PUPJ-DRNA-2022-4-3  

%\bibitem{Bugrov1964} 
%S. Bugrov,
%{\it Approximation of a class of functions with a dominant mixed derivative}. 
%Mat. Sb., 64(106), 410--418, 1964.


% \bibitem{Byrenheid_Dung_Sickel_Ullrich}
% G.\,Byrenheid, D.\,D\~ung, W.\,Sickel, and T.\,Ullrich,
% {\it Sampling on energy-norm based sparse grids for the optimal recovery of {S}obolev type functions in $H^\omega$},
% J. Approx. Theory,
%  207, 207--231,
% 2016,
% https://doi.org/10.1016/j.jat.2016.02.012.

% \bibitem{Byrenheid_Ullrich2017} G. Byrenheid and T. Ullrich,
% {\it Optimal sampling recovery of mixed order Sobolev embeddings
% via discrete Littlewood-Paley type characterizations}, Analisis Matematica, 2017.

% \bibitem{CL08}
% J.-P. Calvi, N. Levenberg, 
% Uniform approximation by discrete least squares polynomials, 
% J. Approx. Theory 152(1), 82--100, 2008.

\bibitem{Carl81}
B. Carl, 
{\it Entropy numbers, s-numbers, and eigenvalue problems},
J. Funct. Anal. 41(3), 290--306, 1981.

\bibitem{Carl-Stephanie-90}
B. Carl and I. Stephani, 
{\it Entropy, compactness and the approximation of operators}, 
Cambridge University Press UK, Cambridge, 1990.

% \bibitem{CD23}
% A. Chkifa and M. Dolbeault, 
% {\it Randomized least-squares with minimal oversampling and
% interpolation in general spaces}, 
% %arXiv:2306.07435.
% SIAM J. Numer. Anal.  62(4), 1515--1538, 2024.
% %DOI: 10.1137/23M160178X

%\bibitem{CKS} F. Cobos, T. Kühn, and  W. Sickel, 
%{\it Optimal approximation of multivariate periodic
%Sobolev functions in the sup-norm},
%Journal of Functional Analysis, 270(11), 4196--4212, 2016.

%\bibitem{CD} A. Cohen, M. Dolbeault, {\it Optimal pointwise sampling for $L_2$ approximation}, Journal of Complexity, 68, 101602, 2022.

% \bibitem{CM} A. Cohen, G. Migliorati, 
% \emph{Optimal weighted least squares methods,} SMAI J. Comput. Math. 3, 181--203, 2017.

\bibitem{Creutzig_Wojtaszczyk2004} 
J. Creutzig, P. Wojtaszczyk,
{\it Linear vs. nonlinear algorithms for linear problems},
J. Complexity 20, 807--820, 2004. 
%doi:10.1016/j.jco.2004.05.003

% \bibitem{DPTT19}
% F. Dai, A. Prymak, V.N. Temlyakov, S.Yu. Tikhonov, 
% {\it Integral norm discretization and related problems}, Russ. Math.
% Surv. 74 (4) (2019) 579--630
% %, arXiv:1807.01353v1
% . Translation from Usp. Mat. Nauk 74:4 (448), 3--58, 2019.

%\bibitem{Diestel-book}
%J. Diestel, 
%{\it Geometry of Banach Spaces -- Selected Topics}, 
%Lecture Notes in Mathematics, Vol. 485. Springer-Verlag, Berlin-New York, 
%1975, xi+282 pp.

\bibitem{DKU} M.\,Dolbeault, D.\,Krieg, M.\,Ullrich,  
{\it A sharp upper bound for sampling numbers in $L_2$}, 
{Appl. Comput. Harmon. Anal.} 63, 113--134, 2023. 
%		\href{https://doi.org/10.1016/j.acha.2022.12.001}{DOI:10.1016/j.acha.2022.12.001}. 
%arXiv:2204.12621, 2022.


% \bibitem{DD91} D.\,D{\~u}ng, {\it On optimal recovery of multivariate periodic functions}, In: Har-
% monic Analysis (Conference Proceedings, Ed. S. Igary), Springer-Verlag
% 1991, Tokyo-Berlin.


%\bibitem{DTU18}
%{D.}~D\~ung, V.\,N.~Temlyakov, and T.~Ullrich,
%\newblock {H}yperbolic {C}ross {A}pproximation.
%\newblock {\em Advanced Courses in Mathematics - CRM Barcelona}.
%  Springer International Publishing, 2018.

\bibitem{Edmunds_Triebel}
D. E. Edmunds, H. Triebel, {\it Function spaces, entropy numbers, differential operators}. Cambridge University Press UK, Cambridge, 1996.

% \bibitem{FHJU22} F. Filbir, R. Hielscher, T. Jahn, and T. Ullrich, {\it Marcinkiewicz--Zygmund inequalities for scattered and random data on the $q$-sphere}, arXiv:2303.00045. 


% \bibitem{Freud} G. Freud,
% {\it Orthogonale Polynome}. Lehrbücher und Monographien aus dem Gebiete
% der Exakten Wissenschaften, Mathematische Reihe, Band 33. Birkh\"auser Verlag, Basel-Stuttgart, 1969.

%\bibitem{Galeev1978} E. M. Galeev,
%{\it Approximation of classes of functions with several bounded
%derivatives by Fourier sums}.
%\newblock {\em Mat. Zametki}, 23, 197--212, 1978; English
%transl. in Mathematical Notes 23, 1978.

% \bibitem{GePe14}
% D.~Geller, I.Z.~Pesensen,
% \emph{Komogorov and linear width of balls in Sobolev spaces on compact manifolds}, 
% Math. Scand. (2014), Vol. 115, No. 1, pp. 96--122.

% \bibitem{GW23}
% J.~Geng and H.~Wang, 
% {\it On the power of standard information for tractability for $L_\infty$ approximation of periodic functions in the worst case setting}, 
% %arXiv:2304.14748.
% J. Complexity, 80, 101790, 2024.
% % https://doi.org/10.1016/j.jco.2023.101790


% \bibitem{Gin}
% E.~Gin\'e, {\it The addition formula for the eigenfunctions of the Laplacian},
% Advances in Mathematics, 18(1):102--107, 1975.


% \bibitem{Griebel_Knapek2000}
% M.\,Griebel and S.\,Knapek, 
% {\it Optimized tensor-product approximation spaces}, 
% Constr. Approx., 16(4), 525--540, 2000, https://doi.org/10.1007/s003650010010.

% \bibitem{Griebel_Knapek2009} 
% M.\,Griebel and S.\,Knapek, 
% {\it Optimized general sparse grid approximation spaces for operator equations},
% Math. Comp.,
% 78(268), 2223--2257, 2009, https://doi.org/10.1090/S0025-5718-09-02248-0.

% \bibitem{Gr19} K. Gröchenig, {\it Sampling, Marcinkiewicz-Zygmund inequalities, approximation, and quadrature rules}, J. Approx. Theory, 257:105455, 2020.

% \bibitem{Heinrich} S. Heinrich, {\it Random approximation in numerical analysis}, in K. D. Bierstedt and et al.,
% editors, Functional Analysis, pages 123–171. Dekker, New York, 1994.

% \bibitem{HL25}
% D. Henrion, J.B. Lasserre. 
% Approximate D-optimal design and equilibrium measure.
% 2025. hal-04688534v2f

\bibitem{HKV16}
A.~Hinrichs, A.~Kolleck, J.~Vyb\'iral, 
{\it Carl’s inequality for quasi-banach spaces}, 
J. Funct. Anal. 271(8), 2293--2307, 2016.

%\bibitem{HKNPU} A. Hinrichs, D. Krieg, E. Novak, J. Prochno and M. Ullrich, 
%{\it Random sections of ellipsoids and the power of random information},
%Transactions of the American Mathematical Society 374(12), 8691-8713, 2021.

\bibitem{HKNPU2}
A.~Hinrichs, D.~Krieg, E.~Novak, J.~Prochno, M.~Ullrich, 
\newblock {\it On the power of random information}, 
\newblock In: F.~J. Hickernell and P.~Kritzer, eds., 
{\em Multivariate Algorithms and Information-Based Complexity}, 43--64, De Gruyter, 
% Berlin/Boston, 
% Berlin, 
2020.

%\bibitem{HKNVa}
%A. Hinrichs, D. Krieg, E. Novak and J. Vybiral, 
%{\it Lower bounds for the error of quadrature formulas for Hilbert spaces},
%Journal of Complexity, 65, 101544, 2021, https://doi.org/10.1016/j.jco.2020.101544.


% \bibitem{HKNV} A.\,Hinrichs, D.\,Krieg, E.\,Novak, and J.\,Vyb\'iral, 
% {\it Lower bounds for integration and recovery in $ L_2$}, J. Complexity, 101662, 2022. % https://doi.org/10.1016/j.jco.2022.101662.

%  \bibitem{HNV} A.\,Hinrichs, E.\,Novak, and J.\,Vyb\'iral, {\it  Linear information versus function evaluations for $L_2$-approximation}, J. Approx. Theory, 153(1),  97-107, 2008.

% \bibitem{HM} S. Hubbert and T.\,M.~Morton,
% {\it $L_p$-error estimates for radial basis function interpolation on the sphere},
% J. Approx. Theory, 129(1):58--77, 2004.


\bibitem{JUV} T. Jahn, T. Ullrich, F. Voigtlaender,
{\it Sampling numbers of smoothness classes via $\ell^1$-minimization},
J. Complexity 79, 101786, 2023.

%\bibitem{KUV} L. K\"ammerer, T. Ullrich, and T. Volkmer, 
%{\it Worst case recovery guarantees for least squares approximation using random samples}, 
%Constr. Approx., 2021, https://doi.org/10.1007/s00365-021-09555-0.
%arXiv: 1911.10111, 2019, to appear in Constructive Approximation.

% \bibitem{KS} M.\,I.~Kadets and M.\,G.~Snobar,
% %{\it Certain functionals on the Minkowski compactum}, Mathematical Notes, 10:694--696, 1971.
% {\it Some functionals over a compact Minkovskii space}, Math. Notes, 10(4), 694--696, 1971.
% % https://link.springer.com/article/10.1007/BF01106467

% \bibitem{Kashin77} 
% B.\,S.\,Kashin, 
% {\it Diameters of some finite-dimensional sets and
% classes of smooth functions}, 
% Izv. Akad. Nauk SSSR Ser. Mat., 41(2): 334--351, 1977; English transl. in Math. USSR-Izv.,
% 11(2), 317--333, 1977,
% https://doi.org/10.1070/IM1977v011n02ABEH001719.

% \bibitem{KKT21}
% B.~Kashin, S.~Konyagin, and V.~Temlyakov, 
% {\it Sampling discretization of the uniform norm},
% Constr. Approx. 57(2), 663--694, 2023.

% \bibitem{Karlin_Studden1966}
% S. Karlin and W.J. Studden,
% {\it Tchebycheff systems: with applications in
% analysis and statistics},
% Wiley Interscience, New York, 1966.
% 

% \bibitem{KKLT} B. Kashin, E. Kosov, I. Limonova, and V. Temlyakov,
% {\it Sampling discretization and related problems},
% J. Complexity, 101653, 2022.
% %https://doi.org/10.1016/j.jco.2022.101653.

% \bibitem{Kol_Lomako_Tikhonov23}
% Y.\,Kolomoitsev, T.\,Lomako, and S.\,Tikhonov, 
% {\it Sparse grid approximation in weighted {W}iener spaces},
% J. Fourier Anal. Appl.,  29 (19), 2023, https://doi.org/10.1007/s00041-023-09994-2.


% \bibitem{Kam19} L. K\"ammerer,
% {\it Multiple lattice rules for multivariate {$L_{\infty}$} approximation in the worst-case setting},
%    {arXiv:1909.02290v1}.
%  %{2019}.
 
%  \bibitem{Kam_Vol19} L. K\"ammerer and T.  Volkmer, 
% {\it Approximation of multivariate periodic functions based on sampling along multiple rank-1 lattices},
%  {J. Approx. Theory},
%   {246}, {1--27},
%   {2019},
% https://doi.org/10.1016/j.jat.2019.05.001.

% \bibitem{KieWo60}
% J.~Kiefer and J.~Wolfowitz.
% \newblock {\it The equivalence of two extremum problems}, 
% \newblock Canadian J. Math. 12, 363--366, 1960.

\bibitem{KT25}
E.D. Kosov, V.N. Temlyakov,
{\it Bounds for the sampling discretization error and their applications to the universal sampling discretization},
J.~Complexity 90, 101958, 2025. 
% https://doi.org/10.1016/j.jco.2025.101958.

% \bibitem{Kosov_Temlyakov23} 
% E.\,D. Kosov and V.\,N. Temlyakov,
% {\it Sampling discretization of the uniform norm and applications},
% %arXiv:2306.14207, 2023.
% J. Math. Anal. Appl. 538 (2), 128431, 2024.
% %https://doi.org/10.1016/j.jmaa.2024.128431.

\bibitem{Krieg-tract}
D.\,Krieg, 
{\it Tractability of sampling recovery on unweighted function classes}, 
Proc. Amer. Math. Soc. Ser. B 11 (2024), 115-125.
% https://doi.org/10.1090/bproc/216

\bibitem{KNU24} 
D. Krieg, E. Novak, M. Ullrich,
{\it On the power of adaption and randomization},
to appear in Forum of Mathematics, Sigma, 
arXiv:2406.07108, 2025.

\bibitem{KPUU-uniform}
D.\,Krieg, K.\,Pozharska, M.\,Ullrich, T.\,Ullrich,
 {\it Sampling projections in the uniform norm}, 
  % {arXiv:2401.02220}.
J. Math. Anal. Appl. 553(2), 129873, 2026.
% https://doi.org/10.1016/j.jmaa.2025.129873.

\bibitem{KPUU-general}
D.\,Krieg, K.\,Pozharska, M.\,Ullrich, T.\,Ullrich,
{\it Sampling recovery in $L_2$ and other norms}, 
to appear in Math. Comp., 
arXiv:2305.07539, 2023.


% \bibitem{KSUW} D. Krieg, P. Siedlecki, M. Ullrich, and H. Wo\'zniakowski,
% {\it Exponential tractability of $L_2$-approximation with function values},
% Adv. Comput. Math. 49, 18, 2023. %https://doi.org/10.1007/s10444-023-10021-7

% \bibitem{KSmani} D. Krieg, M. Sonnleitner,
% {\it Function recovery on manifolds using scattered data}, arXiv:2109.04106, 2021.


\bibitem{KU1} D.\,Krieg, M.\,Ullrich,
{\it Function values are enough for $L_2$-approximation}, 
Found. Comp. Math., 21(4), 1141--1151, 2021.

\bibitem{KU2} D.\,Krieg, M.\,Ullrich,
{\it Function values are enough for $L_2$-approximation: Part~II}, 
J. Complexity 66, 101569, 2021. % https://doi.org/10.1016/j.jco.2021.101569.

% \bibitem{KV} D.\,Krieg and J.\,Vyb\'iral, {\it New lower bounds for the integration of periodic functions}, %arXiv:2302.02639, 2023.
%  J. Fourier Anal. Appl. 29(41), 2023. %https://doi.org/10.1007/s00041-023-10021-7

% \bibitem{Kunsch2018}
%    R.\,Kunsch,
%  {\it Breaking the curse for uniform approximation in {H}ilbert spaces via {M}onte {C}arlo methods},
%   {J. Complexity},
%   {48},  {15--35},
%   {2018},
% https://doi.org/10.1016/j.jco.2018.04.002.


% \bibitem{KSU} T. Kühn, W. Sickel, and T. Ullrich, {\it Approximation of mixed order Sobolev functions on the $d$-torus -- asymptotics, preasymptotics and $d$-dependence}.
% Constructive Approximation, 42, 353--398, 2015.


%\bibitem{KWW08}
%F.\,Kuo, G.\,W.\,Wasilkowski, and H.\,Wo\'zniakowski,
%{\it Multivariate {$L_\infty$} approximation in the worst case setting over reproducing kernel {H}ilbert %spaces},
%Journal of Approximation Theory, 152(2), 135--160, 2008,
%https://doi.org/10.1016/j.jat.2007.11.006.


%\bibitem{KWW}
%F.~Kuo, G.~W.~Wasilkowski, and H.~Wo\'zniakowski,
%{\it On the power of standard information for multivariate approximation
%  in the worst case setting},
%Journal of Approximation Theory, 158, 97--125, 2009.


%\bibitem{LS} Y. T. Lee and H. Sun, {\it Constructing linear-sized spectral sparsification in almost-linear time}, SIAM Journal on Computing, 47(6), 2315-2336, 2018.



% \bibitem{LS} 
% J. Leveslay and X. Sun, {\it Approximation in rough native space by shifts of smooth kernels on spheres}, J. Approx. Theory
% 133, 269--283, 2005.

% \bibitem{Li_Siegel_24}
% Y. Li, J. W. Siegel, 
% {\it Entropy-based convergence rates of greedy algorithms},
% Mathematical Models and Methods in Applied Sciences 34(5), 
% 779--802, 2024. {https://doi.org/10.1142/S0218202524500143}


% \bibitem{Limonova_Malykhin_Temlyakov24}
% I.V. Limonova, Y.V. Malykhin, and V.N. Temlyakov.
% One-sided discretization inequalities and sampling recovery.
% {\it Russian Math. Surveys}  79(3), 515-545, 2024.
% %doi:10.4213/rm10175e.

% \bibitem{LT} I. Limonova and V. Temlyakov, {\it On sampling discretization in $L_2$}, J. Math. Anal. Appl. 515(2), 126457, 2022. 


%\bibitem{LW} W. Lu and H. Wang, {\it On the power of standard information for tractability for $L_2$-approximation in the randomized setting}, Contemporary Mathematics, 3(1), 1--29, 2022.

%\bibitem{MSS} A. Marcus, D. Spielman and N. Srivastava, {\it Interlacing families II: Mixed characteristic polynomials and the Kadison-Singer problem}, Annals of Mathemathics, 327-350, 2015, http://dx.doi.org/10.4007/annals.2015.182.1.8.

% \bibitem{Maiorov1976}
% V.\,E. Maiorov,  {\it The best approximation of classes $W^r_1(I^s)$  in the space $L_\infty(I^s)$}, 
% %Mat. Zametki 19(5), 699--706, 1976.
% Math. Notes
% %of the Academy of Sciences of the USSR 
% 19(5), 420--424, 1976.
% %https://doi.org/10.1007/BF01142563

\bibitem{MUV15}
S.~Mayer, T.~Ullrich, J.~Vyb\'iral, 
{\it Entropy and sampling numbers of classes of ridge functions}, Constr.~Approx.~42, 231--264, 2015. 
% https://doi.org/10.1007/s00365-014-9267-x

%\bibitem{MV92} R. Meise and D. Vogt, {\it Einf\"uhrung in die Funktionalanalysis},
%\newblock Braunschweig; Wiesbaden: Vieweg, 1992 (Vieweg-Studium; 62: Aufbaukurs Mathematik).

% \bibitem{Mhaskar}
% H.\,N. Mhaskar, {\it Weighted quadrature formulas and approximation by zonal function networks on the sphere}, J. Complexity, 22(3):348–370, 2006.

% \bibitem{Mhaskar2}
% H.\,N. Mhaskar, {\it Eignets for function approximation on manifolds},
% Appl. Comput. Harm. Anal.,
% 29(1):63--87, 2010.


%\bibitem{MP} S. Mendelson and A. Pajor, {\it On singular values of matrices with independent rows}, Bernoulli, 12(5), 761-773, 2006.

%\bibitem{Mit} B.\,S.\,Mityagin, {\it Approximation of functions in $L^p$ and $C$ on the torus}.
%Mathematical Notes, 58, 397--414, 1962.

% \bibitem{NSU} N. Nagel, M. Sch\"afer, and T. Ullrich,
% {\it A new upper bound for sampling numbers}, Found. Comp. Math. 22(2), 445-468, 2022.

% \bibitem{Nevai} P. Nevai. 
% \emph{G\'eza Freud, orthogonal polynomials and Christoffel functions. A case study. } J.
% Approx. Theory, 48(1):3--167, 1986.

%\bibitem{NOU2011} S. Nitzan, A. Olevskii and A. Ulanovskii, {\it A few remarks on sampling of signals with small spectrum}, Proceedings of the Steklov Institute of Mathematics, 280(1), 240-247, 2013.

%\bibitem{NOU2016} S. Nitzan, A. Olevskii and A. Ulanovskii, {\it Exponential frames on unbounded sets}, Proceedings of the American Mathematical Society, vol. 144, no 1, 109-118, 2016, http://dx.doi.org/10.1090/proc/12868.

% \bibitem{NNS} V.\,D.\,Nguyen, V.\,K.\,Nguyen, and W.\,Sickel, {\it $s$-numbers of embeddings of weighted
% Wiener algebras}. J. Approx. Theory, 279, Paper No. 105745, 2022.


\bibitem{Novak} E. Novak, 
{\it Deterministic and stochastic error bounds in numerical analysis},
Lecture Notes in Mathematics 1349, Springer-Verlag, 1988.

\bibitem{NoTr06}
E.~Novak, H.~Triebel, 
\emph{Function spaces in Lipschitz domains and
optimal rates of convergence for sampling}, 
Constr. Approx. 23, 325--350, 2006.


%\bibitem{Rauhut_Ward2012} H. Rauhut, R. Ward,
%{\it Sparse Legendre expansions via $\ell^1$-minimization},
% J. Approx. Theory, 164(5), 517--533, 2012.


\bibitem{NW1}
E.~Novak, H.~Wo\'zniakowski,
\newblock {\em Tractability of multivariate problems. {V}olume~{I}: {L}inear
  information}, volume~6 of {\em EMS Tracts in Mathematics}.
\newblock European Mathematical Society (EMS), Z\"urich, 2008.

%\bibitem{NW2}
%E.~Novak and H.~Wo\'zniakowski.
%\newblock {\em Tractability of multivariate problems. {V}olume {II}: {S}tandard
%  information for functionals}, volume~12 of {\em EMS Tracts in Mathematics}.
%\newblock European Mathematical Society (EMS), Z\"urich, 2010.

% \bibitem{NW3} E.\,Novak and H.\,Wo\'zniakowski, {\it Tractability of multivariate problems. Volume~III: Standard information for operators},
% volume~18 of {\em EMS Tracts in Mathematics}.
% \newblock European Mathematical Society (EMS), Z\"urich, 2012.

%\bibitem{NW16}
%E.~Novak and H.~Wo\'zniakowski.
%\newblock {\it Tractability of multivariate problems for standard and linear
%  information in the worst case setting: part I},
%\newblock {Journal of Approximation Theory}, 207, 177--192, 2016.

%\bibitem{Nguyen_Nguyen_Sickel2021}
%	V. D. Nguyen, V. K. Nguyen, W. Sickel,
%{\it $s$-numbers of embeddings of weighted Wiener algebras},
%arXiv: 2011.07663v2, 2021 


% \bibitem{Osipenko_Parfenov1995}
% K.\,Y. Osipenko and  O.\,G.\,Parfenov, 
% {\it Ismagilov type theorems for linear, {G}el'fand and {B}ernstein
%           {$n$}-widths},
%  {J. Complexity},
%  {11(4)}, {474--492},
%  {1995},
% http://dx.doi.org/10.1006/jcom.1995.1025.

% \bibitem{PePe}
% K. B. Petersen and M. S. Pedersen,
% {\it The Matrix Cookbook},
% Technical University of Denmark, Version 20121115, 2012, 
% http://www2.compute.dtu.dk/pubdb/pubs/3274-full.html

\bibitem{Pie07} A. Pietsch. {\it History of Banach spaces and linear operators}, Birkhäuser Boston, MA, 2007.

\bibitem{Pietsch-ideals} A. Pietsch. {\it Operator ideals}, Deutscher Verlag, Wiss., Berlin, 1978.

% \bibitem{Pietsch-s}
% A. Pietsch, {\it s-Numbers of operators in Banach spaces}, 
% Studia Math. 51, 201--223, 1974.


\bibitem{Pinkus85}
A. Pinkus, {\it n-width in approximation theory}, 
Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge / A Series of Modern Surveys in Mathematics (MATHE3, volume 7), Springer Berlin, Heidelberg, 1985. 
%https://doi.org/10.1007/978-3-642-69894-1.


% \bibitem{PU}
% K. Pozharska and T. Ullrich.
% {\it A note on sampling recovery of multivariate functions in the uniform norm}, 
% SIAM J. Numer. Anal., 60(3),  1363--1384, 2022. 
% %https://doi.org/10.1137/21M1410580. (arXiv:2103.11124v2)

\bibitem{PSV24}
J.~Prochno, M.~Sonnleitner, J.~Vyb\'iral, 
{\it Entropy numbers of finite-dimensional Lorentz space embeddings}, 
to appear in Studia Math., 
arXiv:2404.06058, 2024.


% \bibitem{Rud99}
% M. Rudelson, 
% {\it Almost orthogonal submatrices of an orthogonal matrix}, Isr. J. Math. 111, 143--155, 1999.

%\bibitem{S41}
%R. Salem. 
%{\it A new proof of a theorem of Menchoff},
%Duke Mathematical Journal, 8(2), 269--272, 1941.

\bibitem{Schuett84}
C. Schuett, {\it Entropy numbers of diagonal operators between symmetric Banach spaces}, 
J. Approx. Theory 40, 121--128, 1984.

%\bibitem{SU}       
%W. Sickel and T. Ullrich,  
%{\it Smolyak's algorithm, sampling on sparse grids and function spaces        
%of dominating mixed smoothness},        
%East Journal on Approximations, 13(4), 387--425, 2007.  

% \bibitem{Schmidt1944} 
%  E. Schmidt, 
% {\it \"Uber die nebst ihren Ableitungen orthogonalen Polynomensysteme und das zugeh\"orige Extremum},
% Math. Ann., 119, 165--204, 1944, https://doi.org/10.1007/BF01563739.

% \bibitem{Schn14} R. Schneider, {\it Convex bodies: the Brunn-Minkowski theory}, 2nd expanded ed. (English),
% Encyclopedia of Mathematics and its Applications 151. Cambridge University Press, 2014. 
% % (ISBN 978-1-107-60101-7/hbk; 978-1-139-00385-8/ebook). xxii, 736 p. 

% \bibitem{Sh95}
% B.~Shekhtman, 
% {\it Interpolation of individual functions}, 
% Comput. Math. Appl. 30, 191--196, 1995.

% \bibitem{Sm}        
% S. A. Smolyak,       
% {\it Quadrature and interpolation formulas for tensor products       
% of certain classes of functions},     
% Soviet Mathematics Doklady, 4, 240--243, 1963. 

\bibitem{SU23}
M.\,Sonnleitner, M.\,Ullrich, 
{\it On the power of iid information for linear approximation}, 
J. Appl. Numer. Anal., 
 1,  88--126, 2023. 
% https://doi.org/10.30970/ana.2023.1.88.
%arXiv:2310.12740

%\bibitem{Szego1975} 
% G. Szeg\"{o},
%{\it Orthogonal Polynomials},
%4th edition Amer. Math. Soc., Providence, 1975.


%\bibitem{S-book}
%K.-G. Steffens.
%{\it The History of Approximation Theory: From Euler to Bernstein},

%\bibitem{SAT}
%{\it Surveys in Approximation Theory},
%https://www.emis.de/journals/SAT/.

%\bibitem{StChr08}
%A.~Christmann and I.~Steinwart.
%\newblock {\em Support Vector Machines}.
%\newblock Springer, 2008.

%\bibitem{KuSiUl15}
%T.~{K}\"uhn, W.~{S}ickel, and T.~{U}llrich.
%\newblock Approximation of mixed order {S}obolev functions on the $d$-torus:
%asymptotics, preasymptotics and $d$-dependence.
%\newblock {\em Constr. Approx.}, 42:353--398, 2015.

%\bibitem{Lebedev1965} 
%N. N. Lebedev,  
%{\it Special functions and their applications},
%Prentice-Hall, Inc., 1965.

% \bibitem{Steinwart_Scovel2012}
% I. Steinwart and C. Scovel,
% {\it Mercer’s theorem on general domains:
% on the interaction between measures, kernels,
% and {RKHS}s},
% Constr. Approx., 35, 363--417, 2012, 
% https://doi.org/10.1007/s00365-012-9153-3.

\bibitem{Suk86}
A.~G.~Sukharev, 
{\it On the existence of optimal affine methods for approximating linear functionals},
J.~Complexity 2(4), 317--322, 1986. 
% https://doi.org/10.1016/0885-064X(86)90009-9.

%\bibitem{Tem93}
%V.\,N.\,Temlyakov, 
%{\it On approximate recovery of functions with bounded mixed derivative}, J.
% Complexity, 9,  41--59, 1993, https://doi.org/10.1006/jcom.1993.1004. 

%V.~N. Temlyakov.
%\newblock {\em Approximation of periodic functions}, 
%	Computational Mathematics and Analysis Series, 
%	Nova Science Publishers, Inc., Commack, NY, 1993.

%\bibitem{Tem18}
%V.~N. Temlyakov.
%\newblock {\em Multivariate Approximation}, volume~32 of {\em Cambridge
%  Monographs on Applied and Computational Mathematics}.
%\newblock Cambridge University Press, 2018.

\bibitem{T13}
V.\,N.\,Temlyakov, 
{\it An inequality for the entropy numbers and its application}, J. Approx. Theory 173, 110--121, 2013.
% https://
doi.org/10.1016/j.jat.2013.05.003.

\bibitem{T20} V.\,N.\,Temlyakov, 
{\it On optimal recovery in $L_2$},  
J. Complexity, 65, 101545, 2021. 
% https://doi.org/10.1016/j.jco.2020.101545.

\bibitem{Tem25}
V.\,N.\,Temlyakov, 
{\it Sparse sampling recovery by greedy algorithms}, 
IMA Journal of Numerical Analysis, 2025, draf054.
% https://doi.org/10.1093/imanum/draf054

% \bibitem{T23} V.\,N.\,Temlyakov, {\it 
% On universal sampling recovery in the uniform norm}, %arXiv:2307.04017.
% Proc. Steklov Inst. Math. 323, 206–216, 2023. %https://doi.org/10.1134/S0081543823050139

%\bibitem{T18} V. N. Temlyakov, 
%{\it The Marcinkiewicz-type discretization theorems}, 
%Constructive Approximation, 48, 337-369, 2018.

% \bibitem{TU1}
% V.~N.~Temlyakov and T.~Ullrich. 
% {\it Approximation of functions with small mixed smoothness in the uniform norm},
% Journal of Approximation Theory, 277, 105718, 2022.

%\bibitem{TU2} V. N. Temlyakov and T. Ullrich, 
%{\it Bounds on Kolmogorov widths and sampling recovery for
%classes with small mixed smoothness}, Journal of Complexity, 67, 101575, 2021.

%\bibitem{TWW88}
%J.~Traub, G.~Wasilkowski, and H.~Wo\'zniakowski, \emph{Information-Based Complexity}, 
%Acad.~Press, 1988.

% \bibitem{Triebel-Lipschitz}
% H.~Triebel, {\it Function spaces in Lipschitz domains and on Lipschitz manifolds. Characteristic functions as pointwise multipliers}, Revista Matemática Complutense 15.2, 475--524, 2002.

\bibitem{Triebel-entropy}
H.~Triebel, 
{\it Interpolation properties of $\epsilon$-entropy and diameters. 
Geometric characteristics of imbedding for function spaces of
Sobolev-Besov type},
Mathematics of the USSR-Sbornik, Volume 27, Issue 1, 23--
37, 1975.
% https://doi.org/10.1070/SM1975v027n01ABEH002496

%\bibitem{Tr11} J. Tropp, {\it User-friendly tail bounds for sums of random matrices}, Found. Comp. Math., %12(4):389–434, 2011.

%\bibitem{U2020} M. Ullrich, {\it On the worst-case error of least squares algorithms for $L_2$-approximation %with high probability}, J.~Complexity, 60, 2020.

\bibitem{U24} M. Ullrich, 
{\it Inequalities between s-numbers}, 
Adv. Oper. Theory 9, 82 (2024). 
% https://doi.org/10.1007/s43036-024-00386-x

% \bibitem{V08} J. Vyb\'\i ral, 
% {\it Widths of embeddings in function spaces},
% J. Complexity, 24(4), 545--570, 2008.

% \bibitem{XN23}
% Y.~Xu, A.~Narayan, 
% \emph{Randomized weakly admissible meshes}, 
% J.~Approx. Theory 285, 105835, 2023.

% \bibitem{Zeng_Kritzer_Hickernell2009}
% X. Zeng, P. Kritzer, and F. J. Hickernell,
% {\it Spline methods using integration lattices and digital nets},
% Constr. Approx., 30, 529--555, 2009,
% https://doi.org/10.1007/s00365-009-9072-0.



%\bibitem{WW95}         
%G. W. Wasilkowski and H. Wo\'zniakowski,       
%{\it Explicit cost bounds of algorithms for multivariate  tensor product problems},
%Journal of Complexity, 11, 1--56, 1995.

%\bibitem{WW01}
%G.\,W.~Wasilkowski and H.~Wo\'zniakowski. 
%{\it On the power of standard information for weighted approximation},
%Foundations of Computational Mathematics, 1(4), 417-434, 2001.

%\bibitem{WW07} G.W. Wasilkowski and H. Wo\'zniakowski, 
%{\it The power of standard information for multivariate approximation 
%in the randomized setting}, Mathematics of Computation, 76, 965-988, 2007.


\end{thebibliography}

\documentclass[a4paper,12pt]{amsart}

\usepackage[utf8]{inputenc}
\usepackage{amsmath, amsthm, amsfonts, amssymb}
\usepackage[foot]{amsaddr}
\usepackage[hidelinks]{hyperref} 

\newcommand{\norm}[1]{\left\Vert#1\right\Vert}

\newcommand{\eps}{\varepsilon}

\newcommand{\N}{\mathbb{N}}

\newcommand{\R}{\mathbb{R}}

\newcommand{\abs}[1]{\left|#1\right|}

\newcommand{\set}[1]{\Bigl\{#1\Bigr\}}

\newcommand{\co}{\colon}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}

\newcommand{\lin}{\ensuremath{\mathrm{lin}}}

\newtheorem{theorem}{Theorem}

\newtheorem{prop}[theorem]{Proposition}

\theoremstyle{definition}

\title[Sampling vs.~entropy]{
Sampling and entropy numbers \\ in 
the uniform norm
}

\author{Mario Ullrich}

\date{\today}

\address{
Institute of Analysis \& Department of Quantum Computing, 
Johannes Kepler University Linz, Austria}
\email{mario.ullrich@jku.at}

\allowdisplaybreaks

\keywords{
uniform approximation, 
width, 
covering numbers, 
sampling recovery, 
information-based complexity} 
\subjclass[2010]{
41A65; 
41A46, 
41A50 
}

\begin{document}

\begin{abstract}
We prove a sharp bound between sampling numbers and entropy numbers in the uniform norm 
for bounded convex sets of bounded functions. 

\end{abstract}

\maketitle

Let $F$ be a class of 
real-valued functions on a domain $D$
and let $Y$ be some semi-normed space that contains $F$.
The \emph{$n$-th sampling number} of the class $F$ in $Y$ is defined by
\[
g_n(F,Y) \,:=\, 
\inf_{\substack{x_1,\dots,x_n\in {D}\\ 
\phi\co \R^n\to Y
}}\, 
\sup_{f\in F}\, 
\Big\|f - \phi \big(f(x_1),\dots,f(x_n)\big)\Big\|_Y.
\]
This is the \emph{minimal worst case error} that 
can be achieved 
by 
sampling recovery algorithms 
based on at most $n$ function 
evaluations 
at fixed nodes~$x_i$ 
if the error is measured in $Y$. 
We do not consider adaptive choice of $x_i$. 
Here, we only consider $Y=B(D)$, 
i.e., the space of all bounded 
real-valued functions on $D$ 
equipped with the sup-norm $\|f\|_\infty:=\sup_{x\in D}|f(x)|$.

We present an elementary argument 
that shows that 
$g_n(F,B(D))$ 
is bounded from above 
in terms of
the \emph{entropy numbers}, 
defined by
\[
 \eps_n(F,Y) \,:=\, 
 \inf_{y_1, \dots, y_{2^{n}} \in Y} \, \sup_{f\in F} \, \min_{i=1,\dots,2^n} \|f-y_i\|_Y.
\]
In essence, $\eps_n(F,Y)$ represents the smallest $\varepsilon$ such that 
$F$ can be covered by $2^{n}$ balls in $Y$ of radius $\varepsilon$. 
The entropy numbers are closely related to the \emph{covering numbers}, and a
rapid decay indicates that $F$ can be approximated by a relatively small number of elements, implying a low-dimensional structure or effective compressibility.
The $n$-th entropy number can be interpreted as the minimal worst-case error over all algorithms that have access to $n$ 
\emph{binary measurements}, 
where the $n$ bits encode the 
closest point $y_i$, 
and entropy numbers are therefore often used as benchmark for compression. 
Note that 
$\eps_n$ is bounded (or a null sequence) if and only if $F$ is bounded (or totally bounded). 

\medskip

We now state the main result, Theorem~\ref{thm:main}, and we will discuss related bounds below. 
Note that the main advances 
over previous results are its generality, 
as well as better constants and an easier proof.

\begin{theorem}\label{thm:main}
Let $D$ be a set
and $F\subset B:=B(D)$ be convex and 
bounded.
Then, for every 
$n\in\N_0$, 
we have
\[
g_{n}(F,B)  
\;\le\; (n+1)\cdot \eps_n(F,B). 
\]
\end{theorem}


\bigskip
\goodbreak

In the proof below, 
which is very much inspired by \cite[Theorem~12.3.2]{Pietsch-ideals}, 
we will also use the \emph{inner entropy numbers} 
\[
\varphi_n(F,Y) \,:=\, 
\sup_{y_1, \dots, y_{2^{n}+1} \in F}\, \min_{i\neq j}\, \frac12\,\big\|y_i-y_j\big\|_Y
\]
and the basic fact that 
$\varphi_n(F,Y) \le \eps_n(F,Y)\le 2\cdot \varphi_n(F,Y)$, 
see~\cite{Carl-Stephanie-90} 
or~\cite[Chapter~12]{Pietsch-ideals}.
Moreover, we use 
\[
g_n^{0}(F,Y) \,:=\, 
\inf_{\substack{x_1,\dots,x_n\in {D}}}\, 
\sup_{\substack{f,g\in F\co \\ f(x_k)=g(x_k), \\ k=1,\dots,n}}\, 
\big\|f - g\big\|_Y, 
\]
which is called the \emph{(global) diameter of information}, 
see~\cite[Section~4.1]{NW1}.
Again, we have the general relation 
$g_n(F,Y)\le g_n^0(F,Y) \le 2\cdot g_n(F,Y)$, 
and we will use below that in fact $g_n^0(F,B(D))= 2\cdot g_n(F,B(D))$.

\medskip

\begin{proof}[Proof of Theorem~\ref{thm:main}] 
%
We first show that, for 
every $\rho<g_{n}^0(F,B)$, 
we can find 
$f_0,g_0, \ldots, f_n,g_n \in F$ and 
$x_0,\ldots,x_n \in D$ 
such that
$f_k(x_j)=g_k(x_j)$ for $j<k$ and 
$|f_k(x_k)-g_k(x_k)|>\rho$ 
for $k=0,\ldots,n$. 

The proof is by induction. 
Let 
$k\in\{0,\dots,n\}$ 
and assume that $f_j,g_j$ and $x_j$ for $j<k$ have already been found. 
Define 
\[
M_k \,:=\, \set{(f,g)\in F\times F\co f(x_j)=g(x_j) \;\;\text{ for }\; 0\le j< k}. 
\] 
By definition, 
we can choose $f_k,g_k\in F$ 
with 
$(f_k,g_k)\in M_k$ 
and 
\begin{equation*}
\|f_k-g_k\|_\infty \,>\, \rho. 
\end{equation*}  
Hence,  
there is $x_k\in D$ with 
\begin{equation*} 
    |f_k(x_k)-g_k(x_k)| \;>\; \rho.
\end{equation*}
This finishes the induction step. 

\medskip

For 
$\xi=(\xi_0,\dots,\xi_n)\in\{0,1\}^{n+1}$, 
we now define 
\[
h_\xi:=\frac{1}{n+1} \sum_{i=0}^n \left( \xi_i f_i + (1-\xi_i) g_i \right). 
\]
By convexity, $H_n:=\{h_\xi\co \xi\in\{0,1\}^{n+1}\}\subset F$ with $\#H_n=2^{n+1}$.

For $\xi,\xi'\in H_n$, $\xi\neq\xi'$, 
let $k:=\min\{j\co \xi_j\neq\xi'_j\}$.
Then, 
\[\begin{split}
\abs{h_\xi(x_k)-h_{\xi'}(x_k)} 
\,&=\, \frac{1}{n+1} \abs{\sum_{i=k}^{n} ( \xi_i-\xi'_i) f_i(x_k) + (\xi'_i-\xi_i) g_i(x_k) } \\
\,&=\, \frac{1}{n+1} \abs{\xi_k-\xi'_k} \abs{f_k(x_k) - g_k(x_k) } 
\,\ge\,  \frac{\rho}{n+1}.
\end{split}
\]
This implies 
$\norm{h_\xi-h_{\xi'}}_\infty 
\,\ge\,  \frac{\rho}{n+1}$ 
for all $\xi\neq \xi'\in\{0,1\}^{n+1}$ 
and therefore 
\[
\eps_n(F,B) 
\,\ge\, \varphi_n(F,B) 
\,\ge\, \frac{\rho}{2(n+1)}.
\]
Since this holds for all $\rho<g_{n}^0(F,B)$, we obtain 
$g_{n}^0 \le 2(n+1)\cdot \eps_n$.

We finally show that $g_{n}^0=2\cdot g_n$. 
This is well-known, but we recall a simple proof. 
For this note that a 
reconstruction 
$\phi^*(y)\in B(D)$ for information 
$y\in\R^n$ 
is given by 
the function 
\[
\phi^*(y)(x) := 
\frac12\Biggl(\sup_{\substack{f\in F_y}} f(x)
\,+\,\inf_{\substack{f\in F_y}} f(x)\Biggr)
\]
with $F_y:=F_{y,\{x_1,\dots,x_n\}}:=\{f\in F\co f(x_k)=y_k,\; k=1,\dots,n\}$, 
and $\phi^*(y)=0$ whenever $F_y=\varnothing$. 
By boundedness, $\sup$ and $\inf$ exist. 
This is sometimes called a \emph{Chebyshev center}, 
and it is optimal in a worst case sense over~$F$.
%

With this, 
and taking into account that
$|a-b|=2\cdot|a-m|$ with $m=\frac{a+b}{2}$ for all 
$a,b\in\R$, 
we obtain 
for any choice of 
$x_1,\dots,x_n\in D$ 
that 
\[\begin{split}
g_n(F,B)
\,&\le\, 
\sup_{y\in \R^n}\,
\sup_{f\in F_y}\, 
\big\|f - \phi^*(y)\big\|_\infty 
\,=\, \frac12 
\sup_{y\in \R^n}\,
\sup_{f,g\in F_y}\, 
\big\|f - g\big\|_\infty. 
\end{split}
\]
Taking the infimum over all $x_1,\dots,x_n\in D$ proves 
$2\cdot g_n\le g_n^0$. 
The reverse inequality is obvious. 


\end{proof}

\medskip

The convexity in Theorem~\ref{thm:main} 
is needed. This follows, e.g., from results of~\cite{MUV15}, 
see Remark~4.4, that show that 
$\eps_n$ may decay much faster than~$g_n$.
%
However, it can be relaxed under another condition. \\
In fact, if $F=B_X$ is the unit ball of a $p$-Banach space $X$, $0<p\le1$, i.e., we have 
$\|f+g\|_X^p\le\|f\|_X^p+\|g\|_X^p$ for all $f,g\in X$, 
then the proof works with the only difference that 
the $(n+1)$ in the definition of $h_\xi$ must be replaced by 
$(n+1)^{1/p}$. This implies 
the following proposition. 

\begin{prop} 
Let $D$ be a set 
and $F\subset B:=B(D)$ be the unit ball of 
a $p$-Banach space with $0<p\le1$. 
Then, for every 
$n\in\N_0$, 
we have
\[
g_{n}(F,B) \;\le\; (n+1)^{1/p}\cdot \eps_n(F,B).
\]
\end{prop}

\bigskip



\medskip

If $F\subset B(D)$ is convex and symmetric, 
then Theorem~\ref{thm:main} even holds
for the \emph{linear sampling numbers} 
\begin{align*}
g_n^{\lin}(F,Y) \,:=\, 
\inf_{\substack{x_1,\dots,x_n\in {D}\\ g_1,\dots,g_n\in Y}}\, 
\sup_{f\in F}\, 
\Big\|f - \sum_{i=1}^n f(x_i)\, g_i\Big\|_Y
\end{align*}
with $Y=B(D)$, 
see~~\cite{Creutzig_Wojtaszczyk2004} or \cite[Theorem 4.8]{NW1}.  
Without symmetry, we can restrict to affine algorithms, see~\cite{Suk86}. We omit the details. 


Theorem~\ref{thm:main} is sharp up to constants, as can be seen by applying it
to the unit ball of the 
univariate Sobolev space 
$W^s_p=W^s_p([0,1])$, $s\in\N$, 
on the interval equipped with the norm 
$\|f\|_{W^s_p}:=\|f\|_p+\|f^{(s)}\|_p$. \\
The entropy numbers of these spaces 
are known 
since the 1970s to satisfy 
$\eps_n(W^s_p,L_q)\asymp n^{-s}$ 
for all $1\le p,q\le \infty$ and 
$s\in\N$ ($s>1$ for $p=1$). 
We refer to~\cite[Theorem~5.2]{Birman_Solomyak} for the upper 
and to~\cite[Theorem~3]{Triebel-entropy} for the lower bound, respectively.
See also~\cite{Edmunds_Triebel} and~\cite[6.7.8.15]{Pie07} for further results and history. \\
The sampling numbers 
satisfy $g_n(W^s_p,L_q)\asymp g_n^\lin(W^s_p,L_q)\asymp n^{-s+(1/p-1/q)_+}$ 
with $(a)_+ := \max \{a, 0 \}$, again 
for all $1\le p,q\le \infty$ and $s\in\N$ ($s>1$ for $p=1$),  
see~\cite[Section~1.3.11]{Novak} 
as well as~\cite[Theorem~23 \& Corollary~25]{NoTr06}. 
Noting that $L_\infty$ can be replaced by $B$, 
we have 
$g_n(W^s_p,B)\asymp n^{1/p}\cdot \eps_n(W^s_p,B)$. 
Hence, Theorem~\ref{thm:main} gives the correct order for $p=1$. 


%

\medskip
\goodbreak

Let us now discuss some other results on (linear) sampling numbers for general spaces. 
Here, we only consider compact 
$F\subset C(D)$, i.e., the space of continuous functions, where $D$ is a compact topological space. 
We also let $(D,\mu)$ be a Borel probability space, and $L_p:=L_p(D,\mu)$. \\
These are mild assumptions, 
which ensure that the numbers considered here are the same 
no matter if we choose $Y=B(D)$, $Y=C(D)$ or $Y=L_\infty(D)$. 
However, they can sometimes be weakened further, see the corresponding papers. 

First, 
we are only aware of one similar result that bounds sampling by entropy numbers. 
In fact, it was shown recently in~\cite[Remark~2.2]{KT25} that, 
for all convex and symmetric $F\subset C(D)$, $1\le p<\infty$ and $r\neq 1/2$, we have 
\begin{equation} \label{eq:KT25}
g_n(F, L_p) \,\le\, C'\cdot n^{-r'/p} \; 
\left(\sup_{k\le n}\; (k+1)^r\cdot\eps_k(F,B) \right)^{1/p}
\end{equation}
with $r':=\min\{r,1/2\}$ and $C'>0$ only depending on $r\neq1/2$ and $p$. 
%
This result cannot give a bound smaller than 
$n^{-1/(2p)}$, no matter how fast the $\eps_n$ decay.  
In particular, it does not help for $p=\infty$. \\
In this case, 
a similar result to Theorem~\ref{thm:main}, at least for convex and symmetric sets and in an asymptotic sense, 
could also be obtained from a variation 
of~\cite[Theorem~1.2]{KNU24}.  
In fact, it can be shown that 
\begin{equation}
g_{2n-1}^{\rm lin}(F,B) 
\;\le\; 4 n\,\bigg(\prod_{k<n} \eps_k(F,B) \bigg)^{1/n},     
\end{equation}
by replacing, in the proof of~\cite[Theorem~6.4]{KNU24}, the use of Proposition~4.4 by~\cite[12.3.1]{Pietsch-ideals}, i.e., a bound between entropy and Hilbert numbers for convex and symmetric sets. 
See also~\cite{U24} for the used bounds between s-numbers. 

\medskip 

Some other results 
use 
the 
\emph{Kolmogorov width} 
defined by 
\[
d_n(F,Y) \;:=\; \inf_{\substack{V_n\subset Y\\ \dim(V_n)=n}}\sup_{f\in F}\,\inf_{g \in V_n}\|f-g\|_Y.  
\] 
In particular, 
using a simple least squares estimator with suitably chosen points, it has been shown in 
\cite[Theorem~20]{KPUU-uniform} that 
\begin{equation}
\label{eq:KPUU}
g_{4n}^\lin(F,L_p) 
\;\le\; 84\cdot n^{(1/2-1/p)_+}\;d_n(F,B).
\end{equation}
%
The case $p=2$ was obtained earlier in~\cite[Theorem~1.1]{T20}.
For $p=\infty$, the inequality $g_n\le (n+1)\cdot d_n$ is known from~\cite[Proposition~1.2.5]{Novak}. The latter is best possible, and could be improved by~\eqref{eq:KPUU} only by allowing \emph{oversampling}. 
We refer to~\cite[Section~5]{KPUU-uniform} for details. 
Now,~\eqref{eq:KPUU} is also sharp up to constants as can be seen by the same example as above  and 
$d_n(W^s_p,B)\asymp n^{-s+(1/p-1/2)_+}$, 
see~\cite[Theorem~VII.1.1]{Pinkus85} 
and~\cite[6.7.8.13 (as well as 6.2.5.2)]{Pie07} 
for more on this and the interesting history.

More is known in the case of $L_2$-approximation. 
It was shown in~\cite{KPUU-general}, 
based on~\cite{KU1,KU2,DKU}, that, 
for some absolute constant $c\in\N$, 
we have
\[
g_{cn}^{\rm lin}(F,L_2) 
\;\le\; \frac{1}{\sqrt{n}}\sum_{k > n} \frac{ d_k(F,L_2)}{\sqrt{k}}. 
\] 
This shows that 
$g^\lin_n(F,L_2)\asymp d_n(F,L_2)$ 
whenever
$d_n(F,L_2)\asymp n^{-s}$ with $s>1/2$. 
Moreover, the optimal sampling points can be replaced by a (sometimes feasible) random choice if we allow a $\log(n)$-oversampling, see the above papers or the surveys~\cite{HKNPU2,SU23}. 
This has been extended to other $Y$ in~\cite{KPUU-general}, but only under additional assumptions on~$F$. 


Also note that entropy numbers are related to Kolmogorov width, at least if $F=B_X$ is the unit ball of a (quasi-)Banach space $X$. 
In fact, we have
$d_n(B_X,Y) \le n\cdot \eps_n(B_X,Y)$ 
for Banach spaces $X,Y$, see~\cite[Theorem~12.3.2]{Pietsch-ideals}. 
Moreover, there is \emph{Carl's inequality}~\cite{Carl81}, 
i.e., 
$$\eps_n(B_X,Y) \,\le\, c_s\, n^{-s} \cdot \sup_{k\le n}\, k^s\cdot d_k(B_X,Y)$$ with a constant $c_s>0$ only depending on $s>0$, 
which also holds for quasi-Banach spaces, see~\cite{HKV16}, 
and with $d_k$ replaced by Gelfand widths/numbers.
However, it seems that none of the above results implies another.

Let us finally mention that there has been quite some interest in 
bounding sampling numbers by \emph{best-$m$-term approximation}, 
which is another quantity aiming for nonlinear approximation 
like entropy. 
We did not find any interesting corollary from Theorem~\ref{thm:main}  in this respect and leave it for further research. 
Let us only mention a few references 
that showed superiority of nonlinear sampling algorithms~\cite{JUV,Krieg-tract,Tem25}, 
and a possibly related bound on entropy, see~\cite{T13} and~\cite[Proposition~15]{PSV24}.

\bigskip

\noindent\textbf{Acknowledgement.} 
I would like to thank 
Matthieu Dolbeault, David Krieg and Erich Novak 
for helpful comments. 
Especially, I thank Mathias Sonnleitner who observed that 
the bound of Theorem~\ref{thm:main} is sharp up to a factor of $3$. This can be deduced from computations of~\cite{Schuett84}. \\ 
This work is supported by 
the Austrian Research Promotion Agency (FFG) through the project FO999921407 (HDcode) funded by the European Union via NextGenerationEU.

\linespread{.97}


\end{document}